OpenStudy (anonymous):

A container states that it should be stored at a temperature below thirty-five degrees Celsius. Convert the temperature to degrees Fahrenheit. Round to the nearest whole number, if necessary. five degrees Fahrenheit thirty-seven degrees Fahrenheit ninety-five degrees Fahrenheit one hundred twenty-one degrees Fahrenheit

OpenStudy (whpalmer4):

If you have a temperature in degrees C, you can convert that to degrees F by multiplying by 9/5 and adding 32.

OpenStudy (whpalmer4):

-35 C is actually very close to the temperature where the two scales cross over: -40 * 9/5 + 32 = -360/5 + 32 = -72 + 32 = -40

OpenStudy (anonymous):

Thank you c: !

OpenStudy (dan815):

If you have a temperature in degrees F, you can convert that to degrees C by subtracting 32 and multiplying by 5/9.

OpenStudy (dan815):

lol

OpenStudy (whpalmer4):

And here again: -40 = (-40 -32) * 5/9 = -72 * 5 / 9 = -360 / 9 = -40 :-)

OpenStudy (anonymous):

So what is the answer @dan815

OpenStudy (anonymous):

thanks for the help @dan815